### v1

# Shortcuts to adiabaticity for coherent atom transport in an adjustable family of two-dimensional optical lattices

Abstract
Motivated by the compelling need for coherent atom transport in a variety of emerging quantum technologies, we investigate such transport on the example of an adjustable family of two-dimensional optical lattices [L. Tarruell et al., Nature (London) 483, 302 (2012)] that includes square, triangular, honeycomb, dimerized, and checkerboard lattices as its special cases; dynamical optical lattices of this type have already been utilized for the demonstration of topological pumping and the realization of two-qubit quantum gates with neutral atoms. At the outset, we propose the appropriate arrangements of acousto-optic modulators that give rise to a frequency imbalance between counterpropagating laser beams, thus leading to the dynamical-lattice effect in an arbitrary direction in the lattice plane. We subsequently obtain the dynamical-lattice trajectories that enable atom transport using shortcuts to adiabaticity (STA) in the form of inverse engineering based on a dynamical invariant of Lewis-Riesenfeld type. We then quantify the resulting atom dynamics using the transport fidelity computed from the numerical solutions of the relevant time-dependent Schrödinger equation. We do so for various choices of the system parameters and transport directions, finding favorable results for the achievable transport times and robustness of the resulting transport to various experimental imperfections.
I Introduction
Recent developments in the realm of emerging quantum technologies – such as quantum sensing based on atom interferometry Nav ; D (a); Rod ; Prem ; Mart and neutral-atom quantum computing (QC) Jak ; Bri ; Wei (a); Moh ; Nem ; Kie ; Shi – as well as those in the area of quantum-state engineering Shi (2); Sto (a); Haa (1, 2); Nau ; Ste (1, 2); StoNau (1, 2) that facilitates those technologies, have underscored the need for a fast, nearly lossless, neutral-atom transport in optically-trapped ensembles Chia (c); Tor ; Che (b); Ste ; Zha ; Nes ; Lu+ ; Hic ; Din ; Hau ; HauSto ; Nor ; Hwa ; Mor ; Cic ; Ton . As a result, coherent atom transport Qic enabled by moving the confining optical trap – either an optical lattice, an array of optical tweezers, or their combination – has lately been attracting considerable attention, on the experimental- and theoretical sides alike. While the renewed experimental interest in this subject pertains largely to the issues related to increased scalability of neutral-atom QC platform Nor , the rekindled theoretical interest has also been motivated by the availability of powerful STA-based quantum-control protocols that can be utilized to model such transport STA ; STAf ; ContRev .
Coherent single-atom transport in an optically-trapped system entails a final atomic state that approximates - as closely as possible - the initial state in the comoving reference frame (i.e. the rest frame of the moving optical trap), up to an irrelevant global phase. This last requirement, tantamount to demanding minimization - ideally a complete absence - of vibrational excitations at the final point of transport, does not, however, necessarily rule out transient excitations at intermediate times Tor . It is this last aspect that motivates one to design moving-trap trajectories using STA-based control protocols STA ; STAf . Generally speaking, such protocols typically lead to the same final states as their adiabatic counterparts, but often require significantly shorter times, thus alleviating the debilitating effects of noise and decoherence. Owing to the fact that adiabatic control-parameter variations typically render some dynamical properties of the system invariant, inverse-engineering schemes based on Lewis-Riesenfeld invariants (LRIs) Lew (a) constitute the most pervasive subclass of STA quantum-control protocols.
The early theoretical investigations of coherent atom transport were mainly restricted to one-dimensional (1D) systems Tor with purely harmonic confining potentials Che (b), a physical situation that can be treated exactly within the LRI-based STA framework. However, given that such simplifications typically do not apply in realistic systems - where there is either a significant coupling between the longitudinal and transverse degrees of freedom, or the relevant confining potential is anharmonic Zha subsequent theoretical studies have been devoted to more practically relevant, higher-dimensional optical lattices of the optical-conveyor-belt Hau and double-well  HauSto types. Atom transport in such systems is enabled by the dynamical-lattice effect, which is induced by appropriately arranged pairs of acousto-optic modulators (AOMs) that give rise to a frequency imbalance between two counterpropagating laser beams. These studies of atom transport in higher-dimensional optical lattices Sto (b); Hof were carried out using the recently proposed eSTA method Whi (a); Odell (1, 2) - a predominantly analytical modification of the STA approach that allows one to account for the anharmonic character of the underlying potentials. The results obtained using this method are comparable to those corresponding to the quantum-speed limit HauSto .
In this paper, we aim to demonstrate – using the STA approach – the feasibility of time-efficient coherent atom transport in an adjustable family of 2D optical lattices proposed in Tarr ; this family of lattices includes square, triangular, dimerized, honeycomb, and checkerboard lattices as its special cases. Dynamical optical lattices that belong to this family have already been utilized experimentally for demonstrating topological pumping Wal and for realizing two-qubit quantum gates with neutral atoms Kie ; these gates originate from exchange interactions between qubits Heu ; Heu (2); Sto (e); Hay .
We start by proposing specific arrangements of AOMs that enable the dynamical-lattice effect in an arbitrary direction in the relevant lattice plane. We then set out to model the ensuing coherent atom transport by employing STA-type protocols, here based on inverse-engineering techniques involving LRIs. Having obtained the STA dynamical-lattice trajectories, we evaluate the resulting single-atom dynamics by solving the corresponding time-dependent Schrödinger equation (TDSE) in the comoving reference frame numerically using the Fourier split-operator method (FSOM). We then quantify the efficiency of STA-based atomic transport through the corresponding transport fidelity for a broad range of values of the relevant system parameters.
In addition to finding STA solutions to the atom-transport problem at hand, we devote significant attention to analyzing their sensitivity to errors, i.e. their robustness to various experimental imperfections, such as deviations in the control amplitudes, the duration of control, the unavoidable presence of noise, etc.
The remainder of this paper is organized as follows. We start Sec. II by providing the basic background of the atom-transport problem at hand, including the Hamiltonian describing such transport and the essential properties of the relevant family of optical lattices; we then specify concrete arrangements of AOMs that enable atom transport in the system under consideration. Section III is devoted to a short survey of the general STA-related results of relevance for the problem at hand. In the following, Sec. IV, we describe how these general results are utilized for obtaining STA dynamical-lattice trajectories and analyze typical examples thereof. Section V is devoted to the description of our approach for the treatment of atomic dynamics. In Sec. VI, we discuss the obtained results for the atom-transport fidelities. To end with, a short summary of the paper - along with the main conclusions drawn - is provided in Sec. VII. In Appendix A we briefly review the application of the FSOM for solving TDSEs, while Appendix B contains a short introduction into the imaginary-time evolution (ITE) approach for computing ground states of quantum systems.
II System and atom transport
To set the stage for further discussion, we introduce the optical-lattice system to be considered in what follows and the principal objectives of our investigation. In particular, in Sec. II.1 below we briefly formulate the problem of atom transport and the characteristic spatial, temporal, and energy scales. We then describe the main features and parameters characterizing the relevant family of 2D optical lattices (Sec. II.2), followed by a discussion of specific configurations of AOMs that give rise to the dynamical-lattice effect in this system in different directions in the xyx-y plane (Sec. II.3).
II.1 Atom transport in optical lattices
The problem to be discussed in what follows is that of transporting an atom of mass $m$ (e.g. of 40K), located at the time $t=0$ in one of the minima of a 2D optical-lattice potential, to another one of its minima, which is situated at a distant location in the xyx-y plane. In keeping with the conventional atom-transport phenomenology, we assume hereafter that the transport distance $d$ is at least an order of magnitude larger than the linear dimension (i.e. the spatial extent) $l_{0}$ of the atomic ground-state wave packet, that is, $d\gtrsim 10\>l_{0}$. The principal assumption inherent to atom transport in optical lattices is that an atom initially resides in the ground state of the potential well corresponding to a lattice-potential minimum and – having covered the distance $d$ at the time $t=t_{f}$ – also finds itself in the ground state of the potential well at its destination. In other words, the atom transport process amounts ideally to a perfect state transfer – from one ground state to its spatially dislocated counterpart described by the same wave function (up to a physically unobservable global phase) – and is not being accompanied by the creation of any excitations.
It will be assumed in this work that atom transport occurs by way of the dynamical-lattice effect; in the simplest scenario this effect is enabled by the interference of two counterpropagating laser beams with a mutual frequency imbalance. Such an imbalance can straightforwardly be created through the use of appropriately arranged pairs of AOMs (see Sec. II.3 below).
The generic form of a time-dependent Hamiltonian describing single-atom transport in the xyx-y plane of a 2D optical lattice with the potential $U_{\mathrm{L}}(\mathbf{r})\equiv U_{\mathrm{L}}(x,y)$ reads
$H_{\textrm{L}}(t)=--\frac{\hbar^{2}}{2m}\>\left(\frac{\partial^{2}}{\partial x^{2}}+\frac{\partial^{2}}{\partial y^{2}}\right)+U_{\mathrm{L}}\left[\mathbf{r}-\mathbf{q}_{0}(t)\right]\>,$ (1)
where $\mathbf{q}_{0}(t)\equiv\left[q_{0}^{\mathrm{x}}(t),q_{0}^{\mathrm{y}}(t)\right]^{\intercal}$ is the time-dependent 2D vector describing the dynamical-lattice trajectory, i.e. the path of a potential minimum. The TDSE corresponding to this last Hamiltonian is typically discussed in the comoving frame. The wave function $\Phi(\mathbf{r},t)$ of an atom in this last reference frame is obtained from its lab-frame counterpart $\Psi(\mathbf{r},t)$ via a generalized Galilean transformation Got . More precisely, $\Phi(\mathbf{r},t)\equiv\hat{\mathcal{U}}(t)\Psi(\mathbf{r},t)$, where in the Hilbert space of the system at hand the unitary operator $\hat{\mathcal{U}}(t)$ is given by (see, e.g. SLK )
$\hat{\mathcal{U}}(t)=e^{i\hat{\mathbf{p}}\cdot\mathbf{q}_{0}(t)/\hbar}\>e^{-im\hat{\mathbf{r}}\cdot\dot{\mathbf{q}}_{0}(t)/\hbar}\>,$ (2)
with $\hat{\mathbf{r}}$ and $\hat{\mathbf{p}}$ being the single-atom (2D) position and momentum operators, respectively.
Based on the Hamiltonian in Eq. (1), in Sec. IV below we discuss how – using an STA-based method – one can design trajectories that enable fast, coherent atom transport in the 2D optical-lattice system under consideration. In particular, both of our approaches (see Sec. II.3 below) enable atom transport along an arbitrary direction in the xyx-y plane; by way of example, we will mostly be interested in transport along the $x$-, $y$-, and diagonal directions, with the respective transport distances being denoted by $d_{\mathrm{x}}$, $d_{\mathrm{y}}$, and $d_{\mathrm{r}}$.
Before closing these preliminary considerations, it is prudent to specify the characteristic time-, length-, and energy scales in the problem under consideration. In particular, the oscillation periods $T_{\mathrm{x}}\equiv 2\pi/\omega_{\mathrm{x}}$ and $T_{\mathrm{y}}\equiv 2\pi/\omega_{\mathrm{y}}$, corresponding to the harmonic approximation of the full optical potential [cf. Eq. (7) in Sec. II.2 below] in the $x$- and $y$ directions, will serve as the relevant time scales of atom transport in the respective directions. Likewise, the harmonic-oscillator zero-point lengths $l_{\textrm{x}}\equiv\sqrt{\hbar/(2m\omega_{\mathrm{x}})}$ and $l_{\textrm{y}}\equiv\sqrt{\hbar/(2m\omega_{\mathrm{y}})}$ will play the role of the characteristic length scales. Finally, the recoil energy $E_{\textrm{R}}\equiv\hbar^{2}k_{\textrm{L}}^{2}/(2m)$, where $k_{\textrm{L}}$ is the magnitude of the relevant laser wave vector $\mathbf{k}_{\textrm{L}}$, will serve as the characteristic energy scale.
II.2 Adjustable family of 2D optical lattices
In what follows, we describe the setup resulting in an adjustable family of 2D dynamical optical lattices that will be utilized as a test bed for the study of atom transport in the remainder of this work. In order to provide a more concise and clearer explanation of the envisioned setup, we hereafter restrict our discussion to the strong confining potentials in the xyx-y plane; we neglect the minor modulations arising from the ever-present confinement of the system in the $z$ direction, which in the anticipated physical realization is close to being harmonic.
The sought-after adjustable family of 2D optical lattices is constructed from an arrangement of three distinct retro-reflected laser beams of linear polarization and wavelength $\lambda_{\mathrm{L}}$ Tarr ; therefore, the entire system can be parametrized by a single wave number $k_{\mathrm{L}}\equiv 2\pi/\lambda_{\mathrm{L}}$. These three beams will be denoted by $\overline{X}$, $X$, and $Y$ in the following (for an illustration, see Fig. 1), with $I_{\overline{X}}$, $I_{X}$, and $I_{Y}$ being the respective beam intensities.
In the envisioned setup Tarr , beams $\overline{X}$ and $X$ propagate along the same axis, while beam $Y$ is oriented along a $90^{\circ}$-rotated axis (cf. Fig. 1). In particular, the interference of two perpendicular beams – namely, beams $X$ and $Y$ – creates a checkerboard lattice with period $\lambda_{\mathrm{L}}/\sqrt{2}$. At the same time, a third beam, $\overline{X}$ – which is collinear with $X$, but detuned from it by a frequency $\delta$ – gives rise to an additional standing wave with a period of $\lambda_{\mathrm{L}}/2$.
The potential that corresponds to the envisioned adjustable family of 2D optical lattices can be derived starting from the general expression for the light intensity $I(x,y)$ in terms of the electric field $\mathbf{E}(x,y)$ Sho
$I(x,y)=c\epsilon_{0}\>|\mathbf{E}(x,y)|^{2}\>.$ (3)
By making use of this last expression, one finds that the intensity pattern corresponding to the checkerboard lattice is given by
$I_{1}(x,y)=$ - $I_{X}\,\sin^{2}(k_{\mathrm{L}}x)-I_{Y}\,\cos^{2}(k_{\mathrm{L}}y)$
- $2\alpha\,\sqrt{I_{X}I_{Y}}\sin(k_{\mathrm{L}}x)\cos(k_{\mathrm{L}}y)\cos\phi\>,$
where $\phi$ corresponds to the relative phase between the two arms, while $\alpha$ is the visibility introduced to quantify the influence of beam-quality in an experimental setting. The possible sources of experimental imperfections are, for instance, imperfect coherence and mode-mismatching. In what follows, we will restrict our analysis to high-contrast regimes of $\alpha=0.9$, which is readily achievable in experimental realizations Tarr .
Figure 1: Schematic of a setup that enables the adjustable optical-lattice formed from three folded retro-reflected beams. The direct digital synthesizer (DDS) is used to control the frequency change induced by the acousto-optic modulator (AOM), giving rise to the phase $\theta$. The half-wave plates (HWP) can be adjusted in their angular orientation to set the specific intensities wanted for the three laser beams. The red dot indicates the center of the optical-lattice potential.
By once again employing the general expression for laser-beam intensity [cf. Eq. (3)], one finds that the laser-intensity pattern corresponding to beam $\overline{X}$ is given by
$I_{2}(x,y)=-I_{\overline{X}}\,\cos^{2}(k_{\mathrm{L}}x+\theta/2)\>,$ (5)
where the phase $\theta$ determines the relative position of the checkerboard pattern and the 1D standing wave; physically, $\theta$ originates from a detuning between the two laser beams along the $x$ axis, i.e. beams $\overline{X}$ and $X$ (for a schematic illustration, see Fig. 1 below).
From Eqs. (II.2) and (5) it is straightforward to obtain the expression for the total laser-beam intensity $I_{\mathrm{tot}}(x,y)=I_{1}(x,y)+I_{2}(x,y)$. It then immediately follows that the corresponding total optical-lattice potential $U_{\mathrm{L}}(x,y)\>\propto\>I_{\mathrm{tot}}(x,y)$, describing the envisioned family of 2D lattices, in its most general form reads
$U_{\mathrm{L}}(x,y)=$ - $U_{X}\,\sin^{2}(k_{\mathrm{L}}x)-U_{Y}\,\cos^{2}(k_{\mathrm{L}}y)$
- $U_{\overline{X}}\,\cos^{2}(k_{\mathrm{L}}x+\theta/2)$
- $2\alpha\,\sqrt{U_{X}U_{Y}}\sin(k_{\mathrm{L}}x)\cos(k_{\mathrm{L}}y)\cos\phi\>,$
where $U_{X}$, $U_{\overline{X}}$, and $U_{Y}$ are the three single-beam lattice depths (proportional to the laser-beam intensities). This last optical-lattice potential is closely related to the potential utilized in the experimental realization of an adjustable family of 2D optical-lattice structures by Tarruell et al. Tarr . For the sake of simplifying the analysis of coherent atom transport in the present work, we make the following two changes compared to Tarr . Firstly, we displace the $x$ coordinate by the distance $\pi/(2k_{\mathrm{L}})$ [i.e. make the replacement $x\rightarrow x-\pi/(2k_{\mathrm{L}})$]. Secondly, we adopt a somewhat different convention for the phase $\theta$; more specifically yet, we make the replacement $\theta\rightarrow\theta+\pi$.
The general optical-lattice potential in Eq. (II.2) gives rise – by adjusting the relative magnitude of the potential depths $U_{X}$, $U_{\overline{X}}$, and $U_{Y}$ – to several different lattice configurations; their corresponding potentials – as well as their respective periodically-repeating units (i.e. unit cells) – are visualized in Fig. 2. Among them, the following four lattice structures can be realized: square, dimerized, honeycomb, and 1D chains (cf. Fig. 2).
Figure 2: Density plots of the optical potential for the (a) square, (b) dimerized, (c) honeycomb, and (d) 1D-chains, optical lattices, obtainable by adjusting the different lattice depths present in the expression for $U_{\mathrm{L}}(x,y)$ [cf. Eq. (II.2)]. For each lattice structure, the corresponding unit cell is indicated by the dashed white line. The point $(x_{0},y_{0})=(0,0)$, around which the harmonic approximation of the full potential is developed, lies in the center od the unit cell.
Having analyzed the full, anharmonic optical-lattice potential of the system at hand [cf. Eq. (II.2)], it is pertinent to find at this point its harmonic counterpart, i.e. its approximate form, with terms up to quadratic order in $x$ and $y$. This harmonic potential, to be denoted by $V_{\mathrm{L}}(x,y)$, is of utmost importance for our further discussion of atom transport (see Secs. III.2 and IV below).
Given that the adjustable optical-lattice potential in Eq. (II.2) has minima around $(x_{0},y_{0})=(0,0)$ for all previously discussed configurations (cf. Fig. 2), we will approximate the full potential around the origin. We do so assuming that $k_{\mathrm{L}}|x|\ll 1$, as well as $k_{\mathrm{L}}|y|\ll 1$. Based on these assumptions, we straightforwardly obtain
$V_{\mathrm{L}}=-V_{\mathrm{d},0}+ma_{\mathrm{x}}x+\frac{m}{2}\left(\omega_{\mathrm{x}}^{2}x^{2}+\omega_{\mathrm{y}}^{2}y^{2}\right)\>,$ (7)
where $V_{\mathrm{d},0}$ depends on $\theta$ and is given by
$V_{\mathrm{d},0}=U_{Y}+U_{\overline{X}}\cos^{2}\left(\frac{\theta}{2}\right)\>.$ (8)
The squared frequencies in Eq. (7) are given by
$\omega_{\mathrm{x}}^{2}$ $=$ $\frac{2U_{X}k_{\mathrm{L}}^{2}}{m}\left(\frac{U_{\overline{X}}}{U_{X}}\cos\theta-1\right)\>,$ (9)
$\omega_{\mathrm{y}}^{2}$ $=$ $\frac{2U_{Y}k_{\mathrm{L}}^{2}}{m}\>,$
while the expression for the factor $a_{\mathrm{x}}$, which has dimensions of acceleration, reads
$a_{\mathrm{x}}=\frac{U_{\overline{X}}k_{\mathrm{L}}}{m}\left(\sin\theta-2\alpha\>\frac{\sqrt{U_{X}U_{Y}}}{U_{\overline{X}}}\cos\phi\right)\>.$ (10)
Therefore, the approximated potential has a term linear in $x$. While the expansions of the conventional optical-lattice potentials do not contain such terms, these unusual terms – linear in the spatial coordinate – are also present in the case of the double-well optical lattice HauSto .
Another important property of the approximated potential $V_{\mathrm{L}}$ in Eq. (7), of interest for our further discussion, is that – by contrast to the total lattice potential, which is nonseparable in the xyx-y plane (i.e. it cannot be written as a sum of terms that depend on only on spatial coordinate)– this approximated potential is separable; this last property greatly simplifies the analysis of atom transport governed by this approximated potential.
II.3 AOM configurations for atom transport
To facilitate further discussion of atom transport in the family of optical lattices under consideration, we first specify the relevant arrangements of AOMs that enable transport along $x$, $y$, and diagonal ($r$) directions.
To be able to appreciate the specific configurations of AOMs that give rise to the dynamical-lattice effect in the system at hand (see Fig. 3 below), it is helpful to first recall how the use of an AOM (also known as a Bragg cell) gives rise to the dynamical-lattice effect. AOMs lead to a shift of an incoming frequency by applying an ultrasonic frequency onto a crystal, which results in a density modulation along that crystal. Accordingly, the incoming laser light encounters the density modulations effectively playing the role of a grating and, consequently, leading to Bragg diffraction. Thus, the use of the AOM leads to a frequency shift that is equal to the frequency of the ultrasonic sound wave Hua . In an actual optical setup, the AOMs are connected to different channels of the direct-digital-synthesizer (DDS), which allow one to control the frequencies $f_{1}$ and $f_{2}$ applied to the crystals in the AOMs. While one AOM turns the frequency down, the other one ramps it upward by the same amount; this leads to an imbalance $\Delta f=f_{1}-f_{2}=c\left(|\mathbf{k}_{+}|-|\mathbf{k}_{-}|\right)/\left(2\pi\right)$ between the counterpropagating laser beams with wave-numbers $|\mathbf{k}_{+}|$ and $|\mathbf{k}_{-}|$, respectively. This imbalance engenders a dynamical (moving) optical lattice with velocity $v=\pi\Delta f/k_{\mathrm{L}}$, where $k_{\mathrm{L}}=|\mathbf{k}_{\mathrm{L}}|$ is the laser wave number HauSto .
Figure 3: Adapted schematic of a setup that enables atom transport for the adjustable optical-lattice. The direct digital synthesizer (DDS) is used to control the frequency change induced by the acousto-optic modulator (AOM), giving rise to the dynamical-lattice velocity $v$. The red dot indicates the center of the optical-lattice potential.
Having described the general procedure for incorporating AOMs into the system to enable atom transport, we now discuss how this approach can be implemented in the setup for the adjustable potential shown in Fig. 1. Currently, a single AOM is employed to tune the phase shift $\theta$ present in the optical-lattice potential of Eq. II.2. However, it is important to clarify that this particular AOM does not participate in the shifting procedure. To enable controlled lattice translation, each arm of the setup must first be modified so that the beams are not simply reflected by a mirror at the end of each arm. Instead, additional beam splitters and mirrors must be incorporated to separate the incoming and reflected components in each arm, allowing them to be treated independently. An illustration of the modified configuration for the relevant region is provided in Fig. 3. Once the four incoming beams are separated, individual AOMs can be placed in each path to control the frequency of the corresponding beam. In this way, a dynamical optical lattice can be generated in each arm through an induced frequency difference $\Delta f$. The resulting transport directions can then be represented as a superposition of displacements along each axis. For ideal diagonal transport, the frequency differences in both arms must be tuned identically.
III STA: survey of results relevant for atom transport
To set the stage for modeling coherent atom transport in the family of optical lattices under consideration using the STA- and eSTA methods (see Sec. IV below), we first provide a short survey of these methods, with emphasis on results of relevance for atom transport. We start with a short introduction into the subclass of STA methods that employ inverse engineering based on LRIs (Sec. III.1) and then specialize our discussion to the problem of harmonic transport (Sec. III.2); for the sake of simplicity, we provide an overview of these STA results using the notation that corresponds to 1D systems.
III.1 Invariant-based inverse engineering and the Lewis-Leach family of Hamiltonians
Inverse engineering based on LRIs - the most widely used subclass of STA - have by now become one of the preferred approach for modeling coherent atom transport. The underlying transport theory, advanced in Tor , employs a particular family of quadratic-in-momentum invariants Lew (a). According to that theory, the idealized harmonic-trapping case and that of an arbitrary (anharmonic) trapping potential entail somewhat different treatments. To be more specific, the perfect transport in the latter case requires one, in principle, to utilize compensating forces in the comoving reference frame.
A dynamical invariant of a time-dependent Hamiltonian $H(t)$ is any operator $I(t)$ that satisfies the equation
$\frac{\partial}{\partial t}I(t)+\left[H(t),I(t)\right]=0\>.$ (11)
[…]
is the Lewis-Riesenfeld phase STA . Consequently, the general solution of the TDSE for $H(t)$ can in this case succinctly be written as a weighted sum of transport modes, i.e.
$|\Psi(t)\rangle=\sum_{n}C_{n}\>e^{i\theta^{\textrm{LR}}_{n}(t)}|\Phi_{n}(t)\rangle\>,$ (13)
where $C_{n}$ are constants.
The most important family of Hamiltonians employed in modeling atom transport are those of the Lewis-Leach type Lew (b). These Hamiltonians are given by
$H(t)=\frac{p^{2}}{2m}+V(x,t)\>,$ (14)
where - in the most general case - the potential $V(x,t)$ reads
$V(x,t)=-F(t)x+\frac{m}{2}\>\omega^{2}(t)x^{2}+\frac{1}{\rho^{2}(t)}U\left(\frac{x-\alpha(t)}{\rho(t)}\right)\>,$ (15)
where $\rho(t)$, $\alpha(t)$, $\omega(t)$, and $F(t)$ are arbitrary functions of time satisfying certain auxiliary equations Tor . In particular, the functions $\rho(t)$ and $\omega(t)$ ought to satisfy the nonlinear second-order ordinary differential equation
$\ddot{\rho}+\omega^{2}(t)\rho=\frac{\omega^{2}_{0}}{\rho^{3}}\>,$ (16)
which is known as the Ermakov equation Erm , where $\omega_{0}$ is a constant. At the same time, the functions $\alpha(t)$, $\omega(t)$, and $F(t)$ have to obey the condition
$\ddot{\alpha}+\omega^{2}(t)\alpha=\frac{F(t)}{m}\>.$ (17)
The quadratic-in-momentum dynamical invariant corresponding to the Hamiltonian in Eq. (14), up to an immaterial additive constant, has the form Tor
$I(t)$ $=$ $\frac{1}{2m}\big[\rho(p-m\dot{\alpha})-m\dot{\rho}(x-\alpha)]^{2}$
$+$ $\frac{m}{2}\>\omega^{2}_{0}\left(\frac{x-\alpha}{\rho}\right)^{2}+U\left(\frac{x-\alpha}{\rho}\right)\>.$
As it turns out, for a broad class of transport problems one can take $\rho(t)\equiv 1$ and $\omega(t)\equiv\omega_{0}$, effectively rendering Eq. (16) irrelevant; thus, in those cases Eq. (17) remains the only auxiliary equation.
III.2 Harmonic transport: STA treatment
Having reviewed the fundamentals of invariant-based inverse engineering and its application in the special case of the Lewis-Leach-type Hamiltonians [cf. Eq. (15)], it is now pertinent to specialize our discussion to the important case of harmonic transport. In the following, we recapitulate all the relevant general results required to obtain STA transport trajectories using LRI-based inverse engineering.
In transport problems that involve a rigid harmonic oscillator driven by the “transport function” $q_{0}(t)$ – corresponding to $F(t)=m\omega^{2}_{0}q_{0}(t)$ in Eq. (15) – the Eq. (17) adopts the form characteristic of a forced harmonic oscillator and $\alpha(t)$ can be thought of as a classical-particle trajectory $q_{c}(t)$ that satisfies $\ddot{q}_{c}+\omega_{0}^{2}[q_{c}-q_{0}(t)]=0$. The relevant dynamical invariant in the case of harmonic transport – a special case of the general quadratic-in-momentum invariant in Eq. (III.1) – is given by
$I(t)=\frac{1}{2m}\big[p-m\dot{q}_{c}(t)]^{2}\\ +\frac{m}{2}\>\omega^{2}_{0}\left[x-q_{c}(t)\right]^{2}\>.$ (19)
[…]
$\Psi_{n}(x,t)=e^{i\theta^{\textrm{LR}}_{n}(t)}\Phi_{n}(x,t)\equiv e^{i\theta^{\textrm{LR}}_{n}(t)}e^{\frac{i}{\hbar}\>m\dot{q}_{c}(t)x}\phi_{n}(x-q_{c})\>,$ (20)
[…]
$\theta^{\textrm{LR}}_{n}(t)=-i\left(n+\frac{1}{2}\right)\omega_{0}t+\frac{m}{2}\>\int_{0}^{t}\dot{q}_{c}^{2}(t^{\prime})\>dt^{\prime}\>.$ (21)
In order to ensure that $I(t)$ and $H(t)$ are compatible at initial and final times, one sets $[I(t),H(t)]=0$ for $t=0$ and $t=t_{f}$. This translates into the following boundary conditions for the auxiliary function $q_{c}(t)$:
$q_{c}(0)$ $=$ $0\>,\qquad\qquad\>\>q_{c}(t_{f})=d\>,$
$\dot{q}_{c}(0)$ $=$ $\ddot{q}_{c}(0)=0\>,\quad\dot{q}_{c}(t_{f})=\ddot{q}_{c}(t_{f})=0\>.$ (22)
The principal idea in modeling harmonic transport is to first chose a functional form of $q_{c}(t)$ that satisfies these last boundary conditions – typically a polynomial – and then obtain $q_{0}(t)$ through inverse engineering, i.e. from the above forced-harmonic-oscillator equation.
It is important to stress that, in addition to those in Eq. (III.2), one can enforce further boundary conditions on $q_{c}(t)$ – for example, by setting higher-order time derivatives of $q_{c}(t)$ to zero for $t=0,\>t_{f}$. In this manner, one can construct several different auxiliary functions $q_{c}(t)$ for the same transport problem (e.g. polynomials of varying degrees) and compute the corresponding STA transport trajectories $q_{0}(t)$. This non-uniqueness of the resulting trajectory can be exploited to choose a trajectory that satisfies to the largest extent possible certain additional requirements.
IV Dynamical-lattice trajectories
In what follows, we apply the general STA results reviewed in Sec. III to coherent atom transport in the adjustable family of 2D optical lattices under consideration. We first utilize the STA approach to obtain the classical path of the potential minima in a moving lattice. We do so within the framework of a single-atom Hamiltonian with the harmonically approximated potential of Eq. (7).
In the atom-transport problem under consideration, the $x$- and $y$ components of the (time-dependent) dynamical-lattice trajectory vector will be denoted by $q_{0}^{\mathrm{x}}(\bm{\lambda}_{\mathrm{x}};t)$ and $q_{0}^{\mathrm{y}}(\bm{\lambda}_{\mathrm{y}};t)$. Here $\bm{\lambda}=\left(\bm{\lambda}_{\mathrm{x}},\bm{\lambda}_{\mathrm{y}}\right)^{\intercal}$ is the (real-valued) control vector, where $\bm{\lambda}_{\mathrm{x}}\equiv[\lambda^{(1)}_{\mathrm{x}},\ldots,\lambda^{(6)}_{\mathrm{x}}$] and $\bm{\lambda}_{\mathrm{y}}\equiv[\lambda^{(1)}_{\mathrm{y}},\ldots,\lambda^{(6)}_{\mathrm{y}}$] are the reduced control vectors corresponding to the displacements in the $x$- and $y$ directions, respectively. The latter are assumed to satisfy the following conditions:
$q_{0}^{\mathrm{x}}(\bm{\lambda}_{\mathrm{x}};\>jt_{f}/7)$ $=$ $\lambda^{(j)}_{\mathrm{x}}\quad(\>j=1,\ldots,6\>)\>,$
$q_{0}^{\mathrm{y}}(\bm{\lambda}_{\mathrm{y}};\>jt_{f}/7)$ $=$ $\lambda^{(j)}_{\mathrm{y}}\>$. (23)
The STA dynamical-lattice trajectory vector, whose $x$- and $y$ components are a special case of $q_{0}^{\mathrm{x}}(\bm{\lambda}_{\mathrm{x}};t)$ and $q_{0}^{\mathrm{y}}(\bm{\lambda}_{\mathrm{y}};t)$, corresponds to the control vector $\bm{\lambda}_{0}\equiv\left(\bm{\lambda}_{\mathrm{x},0},\bm{\lambda}_{\mathrm{y},0}\right)^{\intercal}$, with components $\lambda^{(j)}_{\mathrm{x},0}$ and $\lambda^{(j)}_{\mathrm{y},0}$ $(j=1,\ldots,6)$. Thus, the STA trajectory is described by $q_{0}^{\mathrm{x}}(\bm{\lambda}_{\mathrm{x},0};\>t)$ and $q_{0}^{\mathrm{y}}(\bm{\lambda}_{\mathrm{y},0};\>t)$.
We now proceed to apply the LRI-based transport theory to the approximate system Hamiltonian
$H_{\textrm{L},0}(t)=-\frac{\hbar^{2}}{2m}\>\left(\frac{\partial^{2}}{\partial x^{2}}+\frac{\partial^{2}}{\partial y^{2}}\right)+V_{\mathrm{L}}(\mathbf{r}-\mathbf{q}_{0}(t))\>,$ (24)
i.e. a single-atom Hamiltonian in which $V(\mathbf{r})\equiv V(x,y)$ is given by the simplified lattice potential $V_{\mathrm{L}}$ of Eq. (7). This approximated potential reads
$V_{\mathrm{L}}\left[\mathbf{r}-\mathbf{q}_{0}(t)\right]=-V_{\mathrm{d,0}}+ma_{\mathrm{x}}\big[x-q_{0}^{\mathrm{x}}(t)\big]$
$+$ $\frac{m}{2}\big\{\omega_{\mathrm{x}}^{2}\big[x-q_{0}^{\mathrm{x}}(t)\big]^{2}+\omega_{\mathrm{y}}^{2}\big[y-q_{0}^{\mathrm{y}}(t)\big]^{2}\big\}\>.$
It is important to stress that – owing to the separability of this last potential – displacements in the $x$- and $y$ directions can be treated independently. Furthermore, upon restricting the above potential to one direction, i.e. making the replacement $V_{\mathrm{L}}\rightarrow V^{\mathrm{x}}_{\mathrm{L}}\left(x-q_{0}^{\mathrm{x}}(t)\right)\>\delta_{x,u}+V^{\mathrm{y}}_{\mathrm{L}}\left(y-q_{0}^{\mathrm{y}}(t)\right)\>\delta_{y,u}+V_{\mathrm{d,0}}$ (where u=x,yu=x,\,y), this potential (up to an additive constant) adopts the characteristic Lewis-Leach form [cf. Eqs. (14) and (15)].
In the problem under consideration, we have a Lewis-Leach-type Hamiltonian with $U=0$ and $\rho(t)\equiv 1$. As it relates to the remaining time-dependent parameters of the general potential in Eq. (15), the following identifications can be made. In the case of atom transport in the $x$ direction, i.e. the potential $V_{\mathrm{L}}^{\mathrm{x}}$, we set $\omega_{\mathrm{x}}(t)=\omega_{\mathrm{x}}$ and
$F_{\mathrm{x}}(t)=m\left[\omega^{2}_{\mathrm{x}}q_{0,x}(t)-a_{\mathrm{x}}\right]\>.$ (26)
Analogously, for atom transport along the $y$ direction, i.e. potential $V_{\mathrm{L}}^{\mathrm{y}}$, we have $\omega_{\mathrm{y}}(t)=\omega_{\mathrm{y}}$ and
$F_{\mathrm{y}}(t)=m\omega^{2}_{\mathrm{y}}q_{0,y}(t)\>.$ (27)
In keeping with the standard practice Tor , we also add to $V_{\mathrm{L}}^{\mathrm{x}}$ and $V_{\mathrm{L}}^{\mathrm{y}}$ the irrelevant time-dependent global terms $m\omega^{2}_{\mathrm{x}}q^{2}_{0,x}(t)$ and $m\omega^{2}_{\mathrm{y}}q^{2}_{0,y}(t)$, respectively, which do not yield any force. Accordingly, the auxiliary equations of the forced-harmonic-oscillator type here adopt the form
$\ddot{q}_{\mathrm{c,x}}(t)+\omega_{\mathrm{x}}^{2}\left[q_{\mathrm{c}}^{\mathrm{x}}(t)-q_{0}^{\mathrm{x}}(t)\right]=-a_{\mathrm{x}}\>,$ (28)
$\ddot{q}_{\mathrm{c,y}}(t)+\omega_{\mathrm{y}}^{2}\left[q_{\mathrm{c}}^{\mathrm{y}}(t)-q_{0}^{\mathrm{y}}(t)\right]=0\>,$ (29)
where $\mathbf{q}_{\mathrm{c}}(t)\equiv\left[q_{\mathrm{c}}^{\mathrm{x}}(t),q_{\mathrm{c}}^{\mathrm{y}}(t),0\right]^{\intercal}$ is the corresponding 2D classical-particle trajectory.
One peculiar aspect of these last auxiliary equations is the presence of $a_{x}$ on the right-hand-side of Eq. (28), resulting from the terms linear in $x$ in the approximated potentials $V_{\mathrm{L}}$ [cf. Eq. (7)]. The presence of $a_{x}$ renders the problem at hand analogous to transport problems involving a constant force (e.g., gravity) Tor . It should be noted that Eq. (28) can be reduced to the form of Eq. (29) by a slight redefinition of $q_{\mathrm{c}}^{\mathrm{x}}(t)$ – namely, by introducing $q_{\mathrm{c}}^{\mathrm{x}}(t)\equiv q_{\mathrm{c}}^{\mathrm{x}}(t)+a_{x}/\omega_{\mathrm{x}}^{2}$. In this manner, the problem of atom transport in the $x$ direction, equally like the one in the $y$-direction, is reduced to a well-known type of atom-transport problems solvable via inverse engineering Tor .
Figure 4: Path of a potential minimum as a function of time, obtained using the STA approach, for transport times $t_{f}$ comparable to (a), and an order of magnitude longer than (b) the internal timescale $T_{\mathrm{x}}$.
The boundary conditions for $q_{0}^{\mathrm{x}}(t)$ and $q_{0}^{\mathrm{y}}(t)$ – along with those for for $q_{\mathrm{c}}^{\mathrm{x}}(t)$ and $q_{\mathrm{c}}^{\mathrm{y}}(t)$ consistent with Eqs. (28) and (29) – motivate one to seek the general solution for the path of the potential minimum in the form of the ninth-degree polynomial Hau
$q_{0,u}(t)=d_{u}\sum_{n=3}^{9}b_{n,u}\left(\frac{t}{t_{\mathrm{f}}}\right)^{n}\quad(u=x,y)\>,$ (30)
where $d_{u}$ is the transport distance. From the relevant initial/boundary conditions Hau , the following results for constants $b_{n,u}$ are obtained:
$b_{3,u}$ $=$ $2520\>(t_{\mathrm{f}}\omega_{u})^{-2}\quad,\quad b_{4,u}=-12600\>(t_{\mathrm{f}}\omega_{u})^{-2}\>,$
$b_{5,u}$ $=$ $22680\>(t_{\mathrm{f}}\omega_{u})^{-2}+126\>,$
$b_{6,u}$ $=$ $-17640\>(t_{\mathrm{f}}\omega_{u})^{-2}-420\>,$ (31)
$b_{7,u}$ $=$ $5040\>(t_{\mathrm{f}}\omega_{u})^{-2}+540\>,$
$b_{8,u}$ $=$ $-315\quad,\quad b_{9,u}=70\>.$
The obtained solution for the STA dynamical-lattice trajectory – more precisely, the $x$ component $q_{0,x}$ of the path of the potential minimum – is shown as a function of the transport time $t_{\mathrm{f}}$ in Fig. 4.
V Calculation of atomic dynamics
In what follows, we briefly present our chosen approach for evaluating the single-atom dynamics corresponding to STA-based dynamical-lattice trajectories [cf. Sec. IV]. This approach is based on the use of the FSOM for numerically solving the relevant TDSE (for a short introduction into this method, see Appendix A).
The crucial prerequisite for computing single-atom dynamics in the problem at hand amounts to evaluating the ground-state wave-function of the relevant optical-lattice potential. Within the general framework of the atom-transport problem (cf. Sec. II.1), that wave function represents the initial ($t=0$) atomic wave packet, i.e. the initial condition for single-atom transport. This wave function can be obtained to requisite accuracy using the plane-wave-expansion method (for details, see, e.g., Sto (b)), which entails Fourier expansions of the underlying lattice potential and the periodic part of single-particle Bloch wave functions in terms of reciprocal-lattice vectors. Yet, the same can be accomplished by employing the FSOM in conjunction with the ITE approach (for details, see Appendix B). Given that in this study we employ the FSOM for the dynamics part, this last scheme appears as the more convenient choice.
Given the character of the problem that we are concerned with, it is convenient to switch from the lab frame to the comoving frame via a generalized Galilean transformation [cf. Eq. (2)]. In this manner, the lab-frame single-atom TDSE
$i\hbar\>\frac{\partial}{\partial t}\Psi(\mathbf{r},t)=\left[\frac{\hat{\mathbf{p}}^{2}}{2m}+U_{\mathrm{L}}(\mathbf{r}-\mathbf{q}_{0}(t))\right]\Psi(\mathbf{r},t)\>,$ (32)
is transformed into its counterpart in the comoving frame
$i\hbar\>\frac{\partial}{\partial t}\Phi(\mathbf{r},t)$ $=$ $\Big[\frac{\hat{\mathbf{p}}^{2}}{2m}+\frac{m}{2}\>\dot{\mathbf{q}}_{0}^{2}(t)+m\ddot{\mathbf{q}}_{0}\cdot\mathbf{q}_{0}$ (33)
$+$ $m\mathbf{r}\cdot\ddot{\mathbf{q}}_{0}(t)+U_{\mathrm{L}}(\mathbf{r})\Big]\Phi(\mathbf{r},t)\>.$
It is important to underscore that the terms $m\dot{\mathbf{q}}_{0}^{2}(t)/2$ and $m\ddot{\mathbf{q}}_{0}(t)\cdot\mathbf{q}_{0}(t)$ in Eq. (33) merely lead to time-dependent global phase factors; the latter are immaterial for the atom-transport problem under consideration and can, accordingly, hereafter be safely neglected. Consequently, the total potential that a transported atom experiences in the comoving frame is given by
$W_{\mathrm{L}}(\mathbf{r},t)\equiv U_{\mathrm{L}}(\mathbf{r})+m\mathbf{r}\cdot\ddot{\mathbf{q}}_{0}(t)\>.$ (34)
[…]
$\Psi(\bm{r},t$ $+$ $\delta t)=\exp\left[-\frac{i}{\hbar}\>U_{\mathrm{L}}(\bm{r})\frac{\delta t}{2}\right]T_{\mathbf{r},\mathbf{p}}(\delta t)$ (35)
$\times$ $\exp\left[-\frac{i}{\hbar}\>U_{\mathrm{L}}(\bm{r})\frac{\delta t}{2}\right]\Psi(\bm{r},t)+\mathcal{O}(\delta t^{3})\>,$
where $\hat{T}_{\mathbf{r},\mathbf{p}}(\delta t)$ stands for the operator
$\hat{T}_{\mathbf{r},\mathbf{p}}(\delta t)\equiv\exp\left[-\frac{i}{\hbar}\left(\frac{\hat{\mathbf{p}}^{2}}{2m}\>\delta t+m\hat{\mathbf{r}}\cdot\int_{t}^{t+\delta t}\ddot{\mathbf{q}}_{0}(t^{\prime})\mathrm{d}t^{\prime}\right)\right]\>.$ (36)
By making use of the Baker-Campbell-Hausdorff formula and its special case known as the Weyl identity Gil , $\hat{T}_{\mathbf{r},\mathbf{p}}(\delta t)$ can be recast as (for a detailed derivation, see Appendix C in HauSto )
$\hat{T}_{\mathbf{r},\mathbf{p}}(\delta t)$ $=$ $\exp\left[-\frac{i}{\hbar}\frac{\hat{\mathbf{p}}^{2}}{2m}\>\delta t\right]\,\exp\left[-\frac{i}{2\hbar}\>\hat{\mathbf{p}}\cdot\delta\dot{\mathbf{q}}_{0}(t)\>\delta t\right]\,$ (37)
$\times$ $\exp\left[-\frac{i}{\hbar}\>m\hat{\mathbf{r}}\cdot\delta\dot{\mathbf{q}}_{0}(t)\right]\>,$
where $\delta\dot{\mathbf{q}}_{0}(t)\equiv\dot{\mathbf{q}}_{0}(t+\delta t)-\dot{\mathbf{q}}_{0}(t)$. It is worthwhile to mention that two additional multiplicative terms – that do not depend on the momentum- and coordinate operators, hence leading to irrelevant time-dependent global phase factors – have been dropped on the RHS of the last equation, analogously to the dropped terms in Eq. (33) above.
With the aid of an analog of Eq. (A) and by performing a Fourier transformation of the $\hat{\mathbf{p}}-$dependent terms in (37), switching also to the coordinate representation of the momentum operator [i.e. $\hat{\mathbf{p}}\rightarrow(\hbar/i)\nabla$], we arrive at the FSOM-based (second-order) time-stepping scheme for the atom-transport problem at hand:
$\Phi(\mathbf{r},t$ $+$ $\delta t)=\exp\left[-\frac{i}{\hbar}\>U_{\mathrm{L}}(\mathbf{r})\frac{\delta t}{2}\right]\mathrm{F}^{-1}\Bigg[e^{-i\frac{\hbar k^{2}}{2m}\delta t}$ (38)
$\times$ $\exp\left[-\frac{i}{2}\>\mathbf{k}\cdot\delta\dot{\mathbf{q}}_{0}\delta t\right]\mathrm{F}\Bigg\{\exp\left[-\frac{i}{\hbar}\>m\mathbf{r}\cdot\delta\dot{\mathbf{q}}_{0}\right]$
$\times$ $\exp\left[-\frac{i}{\hbar}\>U_{\mathrm{L}}(\mathbf{r})\frac{\delta t}{2}\right]\Phi(\mathbf{r},t)\Bigg\}\Bigg]+\mathcal{O}(\delta t^{3})\>.$
In order to alleviate the computational burden in the problem at hand, we exploit the discrete translational symmetry of the system and restrict ourselves to the displacement of one unit cell of the original optical lattice. The appropriate choice of this unit cell has to be consistent with the transport direction. For instance, in order to evaluate transport in the $x$- and $y$ directions we make use of a rectangular unit cell (indicated by white dashed lines in Fig. 2).
Yet another problem-specific feature that allows one to further reduce the level of computational difficulty in the problem at hand is closely related to its description in the comoving frame. Namely, the form of the relevant TDSE [cf. Eq. (33)] immediately implies that there is no need to compute the potential term after each time step, but only the correction term $m\mathbf{r}\cdot\ddot{\mathbf{q}}_{0}$ [cf. Eq. (36)] that depends on the acceleration $\ddot{\mathbf{q}}_{0}$ of the potential minimum.
The spatial Fourier transformation is carried out using the FFT algorithm (cf. Appendix ANRc . To be more specific, we make use of a discrete three-dimensional $N_{x}\times N_{y}\times N_{z}$ grid with $N_{x}=200$, $N_{y}=300$, and $N_{z}=100$ points in the computational window whose dimensions (in units of the respective characteristic lenghtscales $l_{\mathrm{x}}$, $l_{\mathrm{y}}$, and $l_{\mathrm{z}}$) are $100$, 100100, and $500$, respectively. For the propagation of the TDSE in the comoving frame [cf. Eq. (33)], we make use of the adaptive approach NRc with a maximal relative error of $10^{-4}$; the latter could – strictly speaking – require additional (smaller in size) substeps in time. Yet, it turned out to be sufficient to take $N_{t}$ between $20$ and $100$ in all our runs.
VI Results and discussion
In what follows, we quantify and discuss the efficiency of STA-based coherent single-atom transport (cf. Sec. IV above) in various optical lattices belonging to the lattice family under consideration. The main figure of merit used for this purpose is the transport fidelity, defined as
$\mathcal{F}(t_{f})=|\braket{\Psi_{\textrm{target}}|\Psi(t_{f})}|^{2}\>.$ (39)
Its dependence on the transport time $t_{f}$ is determined by the overlap of the target state $|\Psi_{\textrm{target}}\rangle$ (the ground state of the displaced optical-lattice potential, computed using the FSOM combined with the ITE approach; see Appendix B) and the actual final (time-evolved) atomic state $|\Psi(t_{f})\rangle$ (evaluated using FSOM-based numerical propagation of the TDSE in the comoving frame; see Sec. V above and Appendix A).
Figure 5: Dependence of the atom-transport fidelity on the transport time $t_{\mathrm{f}}$ for different transport distances $d_{x}$ and potential depths using the STA scheme. The ratio of the lattice depths corresponds to that of a honeycomb lattice. Panels (a)-(c) show results for potential depths of $(U_{X},U_{\overline{X}},U_{Y})=(200,600,200)E_{R}$, $(400,1200,400)E_{R}$, and $(800,2400,800)E_{R}$, respectively.
Figure 6: Dependence of the atom-transport fidelity on the transport time $t_{\mathrm{f}}$ for different transport distances $d_{x}$ using the STA scheme. The ratio of the lattice depths corresponds to that of a square lattice. Panels (a)-(c) show results for potential depths of $(U_{X},U_{\overline{X}},U_{Y})=(0,200,200)E_{R}$, $(0,400,400)E_{R}$, and $(0,800,800)E_{R}$, respectively.
Figure 7: Dependence of the atom-transport fidelity on the transport time $t_{\mathrm{f}}$ for different transport distances $d_{x}$ using the STA scheme. The ratio of the lattice depths corresponds to that of a linear lattice (1D chains). Panels (a)-(c) show results for potential depths of $(U_{X},U_{\overline{X}},U_{Y})=(50,1000,200)E_{R}$, $(100,2000,400)E_{R}$, and $(200,4000,800)E_{R}$, respectively.
Figure 8: Dependence of the atom-transport fidelity on the transport time $t_{\mathrm{f}}$ for different transport distances $d_{x}$ using the STA scheme. The ratio of the lattice depths corresponds to that of a dimerized lattice. Panels (a)-(c) show results for potential depths of $(U_{X},U_{\overline{X}},U_{Y})=(200,400,200)E_{R}$, $(400,800,400)E_{R}$, and $(800,1600,800)E_{R}$, respectively.
The obtained results for the transport fidelity in the optical lattices of the adjustable family under consideration are depicted in Figs. 5-8. One common denominator of all these results, irrespective of the chosen values of the system parameters (such as the lattice depth, transport distance, beam waists, etc.), is the collapse of transport – manifested by a precipitous decrease in the fidelity – for very short transport times $t_{f}$. This characteristic breakdown of transport originates from the fact that for atomic accelerations above a certain maximal value $|a_{\textrm{max}}|$ that is proportional to the lattice depth Hau the atom being transported cannot be confined by the optical-lattice potential. This happens as a consequence of the disappearance of the minima of the “tilted” lattice potential experienced by the atom in its comoving (non-inertial) reference frame; compared to the potential in the lab frame, the comoving-frame potential has an additional term linear in the transport coordinate Hau . This characteristic behavior for short transport times is a generic property of coherent atom transport in optical lattices, i.e. it is not special for the concrete family of lattices being considered here.
The aforementioned breakdown of transport occurs when the maximal acceleration of an atom during transport - hereafter denoted by $|\tilde{a}_{\textrm{max}}|$ - becomes larger than $|a_{\textrm{max}}|$. It should be borne in mind that the actual value of $|\tilde{a}_{\textrm{max}}|$ depends on the chosen dynamical-lattice trajectory (i.e. the path of the potential minimum); yet, quite generally it holds that $|\tilde{a}_{\textrm{max}}|\propto t_{f}^{-2}$. It immediately follows from this last result that for deeper lattices – i.e., for a larger value of $|a_{\textrm{max}}|$, which is proportional to the lattice depth – the largest atomic acceleration $|\tilde{a}_{\textrm{max}}|$ reached becomes larger than $|a_{\textrm{max}}|$ for sufficiently short total time $t_{f}$. In other words, for deeper lattices the breakdown of transport (i.e., the abrupt decrease of the transport fidelity) takes place for shorter times $t_{f}$.
Our obtained numerical results for the transport fidelity [cf. Eq. (39)] are displayed in Figs. 5-8 for honeycomb, square, linear (1D chain), and dimerized lattices, respectively, and two different transport distances ($d=100\>l_{x}$ and $d=400\>l_{x}$). These results are consistent with the above general features of coherent atom transport in optical lattices. In other words, they clearly manifest breakdown of transport at short transport times. On the other hand, in the opposite limit of very long transport times $t_{\mathrm{f}}$ the results coincide with the well-known adiabatic limit, i.e. essentially perfect atom transport ($\mathcal{F}\approx 1$) for $t_{\mathrm{f}}\gg\omega_{\mathrm{j}}^{-1}$ ($j=x,y$) is achieved.
In particular, in the case of the honeycomb lattice, with the parameter values chosen as in Fig. 5, the respective transport times $t_{f}$ corresponding to the onset of the breakdown of atom transport are in the range between $7\>T_{\mathrm{x}}$ and $8\>T_{\mathrm{x}}$ for the transport distance $d=100\>l_{x}$, as well as in the range between $16\>T_{\mathrm{x}}$ and $17.5\>T_{\mathrm{x}}$ for the transport distance $d=400\>l_{x}$.
In the case of the square lattice, which corresponds to Fig. 6, the breakdown of atom transport takes place at times in the range between $5.5\>T_{\mathrm{x}}$ and $6\>T_{\mathrm{x}}$ for the transport distance $d=100\>l_{x}$, and between $10\>T_{\mathrm{x}}$ and $12\>T_{\mathrm{x}}$ for $d=400\>l_{x}$. These results are clearly in agreement with the anticipated trend of decreasing with the increase of the lattice depth.
For the linear (1D-chain) lattice that corresponds to Fig. 7, the transport breakdown takes place when $t_{f}$ is between $5.5\>T_{\mathrm{x}}$ and $6\>T_{\mathrm{x}}$ for $d=100\>l_{x}$, and between $10\>T_{\mathrm{x}}$ and $12\>T_{\mathrm{x}}$ for $d=400\>l_{x}$.
Finally, in the case of the dimerized lattice with the parameter values chosen as in Fig. 8, the onset of the transport breakdown occurs for $t_{f}$ between $10\>T_{\mathrm{x}}$ and $13\>T_{\mathrm{x}}$ for the transport distance $d=100\>l_{x}$, and between $14\>T_{\mathrm{x}}$ and $18\>T_{\mathrm{x}}$ for $d=400\>l_{x}$.
VII Summary and Conclusions
To conclude, in this paper we investigated coherent atom transport in an adjustable family of two-dimensional optical lattices. We first specified concrete arrangements of acousto-optic modulators that give rise to the moving-lattice effect in these systems in an arbitrary direction. We subsequently designed the corresponding dynamical-lattice trajectories using shortcuts to adiabaticity (STA). Through numerical propagation of the relevant time-dependent Schrödinger equation – using the previously computed STA-based dynamical-lattice trajectories – we calculated the efficiency ofhe resulting coherent atom transport as described by the atom-transport fidelity.
We demonstrated that the STA approach enables fast atom transport in the considered family of optical lattices, for a wide range of lattice depths and transport directions. In addition, we analyzed the implications of the atom-transport processes for neutral-atom quantum computing based on collisional two-qubit entangling gates.
By contrast to previous theoretical investigations of coherent atom transport, based mostly on simplified scenarios, in the present work we addressed this phenomenon based on the full, anharmonic optical-lattice potential. Thus, the obtained results can be corroborated in future transport experiments in such lattices. Furthermore, our study also complements well the recent works on quantum-state control of Bose-Einstein condensates in optical lattices Amr . Finally, this work may provide the impetus for further experimental implementations of STA-based control schemes for coherent atom transport in optically-trapped neutral-atom systems, the first one of which has been undertaken quite recently Hwa .
Acknowledgements.
V.M.S. is grateful to K. Viebahn for calling his attention to the dynamical optical-lattice family of Tarr . This research was supported by the Deutsche Forschungsgemeinschaft (DFG) – SFB 1119 – 236615297.
Appendix A FSOM-based numerical propagation of TDSEs
In the following, we briefly review the essential ingredients of the use of the FSOM for solving TDSEs.
Let us consider Cauchy-type initial-value problems $\partial_{t}f(\mathbf{r},t)=\hat{A}(\mathbf{r},t)f(\mathbf{r},t)$, where $\hat{A}(\mathbf{r},t)$ is an operator and the initial condition is given by $f(\mathbf{r},t)=f_{0}(\mathbf{r})$. The FSOM is the method of choice for solving such problems in cases where $\hat{A}$ can be written as a sum of an operator that can straightforwardly be diagonalized in real space and another one that can be diagonalized in Fourier space. One of the most common examples of such problems is the TDSE Fei ; Ban . The crucial idea facilitating the use of FSOM for solving TDSEs is to approximate the time-evolution operator of the system under consideration by a product of operators that are either diagonal in real space or in Fourier space.
Consider the TDSE that describes the motion of a particle of mass $m$ in the potential $U(\bm{r})$. The underlying single-particle Hamiltonian is given by
$H=-\frac{\hbar^{2}\nabla^{2}}{2m}+U(\bm{r})\>.$ (40)
To approximate the time-evolution operator that corresponds to this last Hamiltonian we invoke the identity (known as the symmetrized Trotter decomposition) Gil
$e^{i(A+B)\delta t}=e^{iA\delta t/2}e^{iB\delta t}e^{iA\delta t/2}+\mathcal{O}(\delta t^{3})\>,$ (41)
which holds for arbitrary Hermitian operators $A$ and $B$. In the special case $A=U(\bm{r})$, $B=-\hbar^{2}\nabla^{2}/(2m)$ of this last identity, we obtain an accurate second-order time-stepping scheme  Pec
$\Psi(\bm{r},t$ $+$ $\delta t)=\exp\left[-\frac{i}{\hbar}\>U(\bm{r})\frac{\delta t}{2}\right]\exp\left(i\frac{\hbar\nabla^{2}}{2m}\>\delta t\right)$ (42)
$\times$ $\exp\left[-\frac{i}{\hbar}\>U(\bm{r})\frac{\delta t}{2}\right]\Psi(\bm{r},t)+\mathcal{O}(\delta t^{3})$
for the propagation of the wave-function $\Psi(\bm{r},t)$ of a system described by the Hamiltonian in Eq. (40). This approach to evaluating the time evolution – in which the potential part of the total Hamiltonian is applied for a half time step before and after the kinetic part – is often referred to as symmetric splitting (or, alternatively, as Strang splitting Spe ) and its principal advantage compared to the conventional approach is that it reduces the error to $\mathcal{O}(\delta t^{3})$.
The form of Eq. (42) permits independent treatment of the different exponential terms. Therefore, it is pertinent to Fourier-transform the kinetic term, recasting at the same time the RHS of Eq. (42) using the identity
$\exp\left(i\frac{\hbar\nabla^{2}}{2m}\>\delta t\right)\exp\left[-\frac{i}{\hbar}\>U(\bm{r})\frac{\delta t}{2}\right]\Psi(\bm{r},t)$
$=$ $F^{-1}\left[e^{-i\frac{\hbar k^{2}}{2m}\delta t}\>F\left\{\exp\left[-\frac{i}{\hbar}\>U(\bm{r})\frac{\delta t}{2}\right]\Psi(\bm{r},t)\right\}\right]\>,$
where $F[\ldots]$ stands for the Fourier transform of the argument and $F^{-1}[\ldots]$ for its inverse.
By successively applying the time-propagation scheme based on Eq. (42) $N_{t}$ times on an initial wave-function $\Psi(\bm{r},t)$ one obtains a numerical solution of a TDSE at time $t^{\prime}=t+N_{t}\delta t$. In actual numerical implementations of the FSOM, the wave-function is represented on a discrete rectangular spatial grid with $N_{s}$ points and the exact (continuous) Fourier transform is approximated by its discrete counterpart. The computational burden involved in propagating the wave function is dominated by the need to compute its spatial Fourier transformation and its inverse. If carried out using the fast Fourier transform (FFT) algorithm NRc , an elementary step in these transformations entails $\mathcal{O}(N_{s}\log_{2}N_{s})$ operations.
Appendix B Ground-state calculation via the ITE approach
Here we provide a short introduction into the ITE approach, which established itself as one of the most robust computational methods for obtaining ground states of quantum systems And ; Fei . Its relationship to the FSOMa - its counterpart utilized for computing dynamics - bears close analogy to the connection between Lanczos-type diagonalizations and Chebyshev-propagator methods Stojanović and Salom (2019). It is prudent to start our review of the ITE approach using the abstract, representation-independent notation for quantum states.
Consider the problem of obtaining the ground state of a Hamiltonian denoted by $H$ in the following. One starts by expanding a quantum state $|\phi\rangle$ in terms of the eigenstates $\{|\psi_{n}\rangle,\>n=0,1,\ldots\}$ of $H$, i.e. $|\phi\rangle=\sum_{n}c_{n}|\psi_{n}\rangle$, assuming that the sought-after ground state $|\psi_{0}\rangle$ has a nonvanishing contribution to this expansion (i.e. $c_{0}\equiv\langle\psi_{0}|\phi\rangle\neq 0$). Assuming that $|\phi\rangle$ is the initial ($t=0$) state of the system, its states $|\phi(t)\rangle\equiv e^{-iHt/\hbar}|\phi\rangle$ at later times $t>0$ are given by the linear combinations
$|\phi(t)\rangle=\sum_{n}c_{n}\>e^{-iE_{n}t/\hbar}|\psi_{n}\rangle$ (44)
of the time-evolved eigenstates $e^{-iE_{n}t/\hbar}|\psi_{n}\rangle$ of $H$, where $E_{n}$ is the eigenvalue of $H$ that corresponds to the eigenstate $|\psi_{n}\rangle$.
By performing a Wick rotation into the complex plane (i.e. by making the substitution $t\rightarrow\tau=it$), Eq. (44) goes over into
$|\phi(\tau)\rangle=\sum_{n}c_{n}\>e^{-E_{n}\tau/\hbar}|\psi_{n}\rangle\>.$ (45)
Therefore, $|\phi(\tau)\rangle$ is an exponentially-decaying superposition of the eigenstates $|\psi_{n}\rangle$ with the decay rates given by the energy eigenvalues $E_{n}$. Given that the decay rate of the ground state is the smallest one, in the large-$\tau$ limit one arrives at
$|\phi(\tau)\rangle\approx c_{0}\>e^{-E_{0}\tau/\hbar}|\psi_{0}\rangle\>.$ (46)
Therefore, given a trial state $|\phi(\tau=0)\rangle\equiv|\phi\rangle$ with a nonvanishing overlap with the sought-after ground state, the ITE approach yields a monotonously decreasing functional of $\tau$ that converges to the ground-state energy $E_{0}$ of the system as $\tau\rightarrow\infty$, i.e.
$E_{0}=\lim_{\tau\rightarrow\infty}\langle\phi(\tau)|\>H\>|\phi(\tau)\rangle\>.$ (47)
In the atom-transport problem under consideration (cf. Sec. II.1) – and many other problems – one makes use of the coordinate representation, in which a quantum state $|\psi\rangle$ is represented by the wave function $\psi(\mathbf{r})\equiv\langle\mathbf{r}|\psi\rangle$ (with $|\mathbf{r}\rangle$ being the eigenstates of the position operator). The counterpart of Eq. (45) in this representation reads
$\phi(\bm{r},\tau)=\sum_{n}c_{n}e^{-E_{n}\tau/\hbar}\psi_{n}(\bm{r})\>,$ (48)
where $\phi(\bm{r},\tau)\equiv\langle\mathbf{r}|\phi(\tau)\rangle$ and $\psi_{n}(\bm{r})\equiv\langle\mathbf{r}|\psi_{n}\rangle$ are the relevant wave functions. The remaining steps in the application of the ITE approach within this representation are carried out in complete analogy with the above representation-independent expressions (46) and (47).